OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through the given points. (-1,4) and (1,2)

OpenStudy (anonymous):

First you need to determine the slope. m= y2-y1\[m = y1-y2 \div x1-x2\]

OpenStudy (anonymous):

Point-slope form is \[y - y_0 = m(x-x_0)\] Pick 1 point as (x0,y0) and calculate m as @nelsonjedi said above.

OpenStudy (anonymous):

Recheck your slope.

OpenStudy (anonymous):

The slope is -1 and the equation is y-4=-1(x+1).